A body weighs 160N in air and 120N when wholly immersed in a liquid of relative density 0.8. The relative density value of the body is (g=10ms-2)
(a)2.2
(b)3.2
(c)1.2
(d)3.8
1
Expert's answer
2018-11-16T10:37:09-0500

The body equilibrium condition

F=mg-F_b

F=W-gρ_L V

So

V=(W-F)/(gρ_L )=(160-120)/(10×0.8)=5

The relative density of the body

ρ_b=m/V=W/gV=160/(10×5)=3.2